LOW UNCERTAINTY MEASUREMENTS OF LINE PARAMETERS USING FREQUENCY-STABILIZED CAVITY RING-DOWN SPECTROSCOPY

\underline{J. T. HODGES}, National Institute of Standards and Technology, Gaithersburg, MD 20899; D. LISAK, Nicolaus Copernicus University, Torun, Poland; D. J. ROBICHAUD and M. OKUMURA, Division of Chemistry and Chemical Engineering, California Institute of Technology, Pasadena CA 91125; C. E. MILLER, Jet Propulsion Laboratory, Pasadena, CA 91109.

We discuss how the frequency-stabilized cavity ring-down spectroscopy (FS-CRDS) method enables high-precision and high-sensitivity measurements of shapes, intensities, as well as broadening and shifting parameters of absorption lines^{a}. We illustrate the method’s high spectral resolution by observation of sub-MHz wide Doppler-free saturation features in blended near-ir H_{2}O spectra, and we present measurements and models of H_{2}O and O_{2} spectra that yield line parameters with sub%-level relative uncertainties.

aD. Lisak, J. T. Hodges and R. Ciurylo Phys. Rev. A. 73, 012507 2006.